A Bibliography of Publications about SGML, the Standard Generalized Markup Language: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.13

Title word cross-reference

3 [KR12].

5 [VN10].

access [CH10]. analysis [KD10]. App [Sta10]. application [DWYB10]. approach [LWXZ10]. apps [Sta10]. Automata [SSSS10, TS10].

based [DWYB10, HR10]. becomes [MRA10]. Building [Sta10]. business [KD10].

C [Sta10]. calculus [AH10]. Chain [MNS10]. characterization [DWYB10]. closure [TS10]. clustering [HR10, LC10, TG10]. Cocoa [Sta10]. Collections [SKG^{+}10, CMRV10]. common [HR10]. Complexity [MNS10]. composition [AH10]. computing [Poz13]. concise [BNSV10]. consistency [LWHK10]. consistently [LWHK10]. cookbook [Sch10]. core [DWYB10]. CSS [Sch10, Sta10].

data [BPT10, FMW10, RT10]. databases [BCH^{+}10]. Decision [MNS10]. DescribeX [CMRV10]. description [FMW10]. distributed [BPT10]. document [KD10]. documents [HR10, TG10]. Domotic [MRA10]. DTDs [BNSV10]. Dynamic [CKM10].

Efficient [GLWY10]. electronic [KD10]. eliminate [LWHK10]. enabling [LWHK10]. ensure [CH10]. entity [LWXZ10]. Evaluating [SSSS10]. evaluation [GLWY10]. Exploring [CMRV10]. Expressions [MNS10, SSSS10, BNSV10]. extended [CH10].

feature [AH10].

generating [LHC10]. graph [BPT10]. graph-structured [BPT10].

heterogeneous [BCH^{+}10]. HTML [Sta10, VN10]. HTML5 [Pil10].

Improving [LHC10]. incompatibility [MRA10]. Inference [BNSV10, LC10]. information [CH10]. informative [LHC10]. iPhone [Sta10].

JavaScript [Sta10].

keyword [LWHK10, LWXZ10, LC10].

Labeling [CKM10]. languages [FMW10]. Large [SKG^{+}10]. Linear [SSSS10]. logic [TS10].

Making [Sta10]. mapping [BCH^{+}10]. Matching [SSSS10]. materialized [GLWY10]. MathML [KR12]. mining [RT10]. model [CH10]. multi [DWYB10]. multi-core [DWYB10]. multi-thread [DWYB10].

nested [TS10]. networking [DWYB10]. next [FMW10]. novel [LWXZ10].

Objective [Sta10]. Objective-C [Sta10]. OpenMath [KR12]. Optimizing [BCCY10]. oriented [DWYB10].

P2P [BCH^{+}10]. Pattern [SSSS10]. Pattern-Matching [SSSS10].

Performance [DWYB10]. placement [BPT10]. plan [GLWY10]. Problems [MNS10]. processors [DWYB10]. Publications [SKG^{+}10].

query [BCH^{+}10, GLWY10, LWXZ10].

recursive [BCCY10]. Regular [MNS10, BNSV10]. relations [BCCY10]. Restandardize [VN10]. result [LHC10, LC10]. results [LWHK10]. Return [LC10]. rewriting [GLWY10]. running [Pil10].

Schema [BCH^{+}10]. Schemas [MNS10]. Scientific [SKG^{+}10, Poz13]. search [LWHK10, LHC10, LC10]. secure [CH10]. Semantic [TG10]. Semantics [KR12]. service [BPT10]. services [CH10]. snippets [LHC10]. specification [LC10]. spurious [LWHK10]. standards [KD10]. Store [Sta10]. Structural [LWHK10]. structure [HR10]. structured [BPT10]. subtree [LWXZ10]. survey [KD10]. systems [DWYB10].

technique [HR10]. technologies [MRA10]. thread [DWYB10]. Transforming [SKG^{+}10]. Transitive [TS10]. translation [BCH^{+}10]. transparent [MRA10]. tree [BPT10, TS10]. tree-structured [BPT10]. Trees [CKM10].

uniform [AH10]. updates [BCCY10]. user [MRA10]. using [LWXZ10]. utilizing [LHC10].

view [GLWY10]. views [BCCY10].

walking [TS10]. Web [CH10, CMRV10, VN10]. weighted [HR10]. Will [VN10]. without [Sta10].

XACML [CH10]. XML [BCH^{+}10, BCCY10, CKM10, CMRV10, DWYB10, GLWY10, HR10, LWHK10, LWXZ10, LHC10, LC10, MNS10, Poz13, RT10, SKG^{+}10, TG10]. XPath [SSSS10, TS10].

References

Apel:2010:CUF

[AH10] Sven Apel and Delesley Hutchins.

A calculus for uniform feature composition. ACM Transactions on Programming Languages and Systems, 32(5):19:1–19:??, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bramandia:2010:OUR

[BCCY10] Ramadhana Bramandia, Jiefeng Cheng, Byron Choi, and Jeffrey Xu Yu. Optimizing updates of recursive XML views of relations. VLDB Journal: Very Large Data Bases, 18(6):1313–1333, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bonifati:2010:SMQ

[BCH^{+}10] Angela Bonifati, Elaine Chang, Terence Ho, Laks V. Lakshmanan, Rachel Pottinger, and Yongik Chung. Schema mapping and query translation in heterogeneous P2P XML databases. VLDB Journal: Very Large Data Bases, 19(2):231–256, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bex:2010:ICR

[BNSV10] Geert Jan Bex, Frank Neven, Thomas Schwentick, and Stijn Vansummeren. Inference of concise regular expressions and DTDs. ACM Transactions on Database Systems, 35(2):11:1– 11:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Buehrer:2010:DPS

[BPT10] Gregory Buehrer, Srinivasan Parthasarathy, and Shirish Tatikonda. A distributed placement service for graph-structured and treestructured data. ACM SIGPLAN Notices, 45(5):355–356, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chou:2010:EXM

[CH10] Shih-Chien Chou and Chun-Hao Huang. An extended XACML model to ensure secure information access for Web services. The Journal of Systems and Software, 83(1):77–84, January 2010. CO-DEN JSSODM. ISSN 0164-1212.

Cohen:2010:LDX

[CKM10] Edith Cohen, Haim Kaplan, and Tova Milo. Labeling dynamic XML trees. SIAM Journal on Computing, 39(5):2048–2074, ???? 2010. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Consens:2010:EXW

[CMRV10] Mariano P. Consens, Renée J. Miller, Flavio Rizzolo, and Alejandro A. Vaisman. Exploring XML Web collections with DescribeX. ACM Transactions on the Web (TWEB), 4(3):11:1– 11:??, July 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ding:2010:PCM

[DWYB10] Jason Jianxun Ding, Abdul Waheed, Jingnan Yao, and Laxmi N. Bhuyan. Performance characterization of multi-thread and multicore processors based XML application oriented networking systems. Journal of Parallel and Distributed Computing, 70(5):584– 597, May 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fisher:2010:NDD

[FMW10] Kathleen Fisher, Yitzhak Mandelbaum, and David Walker. The next 700 data description languages. Journal of the ACM, 57 (2):10:1–10:??, January 2010. CO-DEN JACOAH. ISSN 0004-5411.

Gao:2010:EEQ

[GLWY10] Jun Gao, Jiaheng Lu, Tengjiao Wang, and Dongqing Yang. Efficient evaluation of query rewriting plan over materialized XML view. The Journal of Systems and Software, 83(6):1029–1038, June 2010. CODEN JSSODM. ISSN 0164-1212.

Hwang:2010:WCS

[HR10] Jeong Hee Hwang and Keun Ho Ryu. A weighted common structure based clustering technique for XML documents. The Journal of Systems and Software, 83(7): 1267–1274, July 2010. CODEN JSSODM. ISSN 0164-1212.

Kabak:2010:SAE

[KD10] Yildiray Kabak and Asuman Dogac. A survey and analysis of electronic business document standards. ACM Computing Surveys, 42(3):11:1–11:??, March 2010. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Kohlhase:2012:SOM

[KR12] Michael Kohlhase and Florian Rabe. Semantics of OpenMath and MathML 3. Mathematics in Computer Science, 6(3):235– 260, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=1661-8270&volume=6&issue=
                                        3&spage=235.

Liu:2010:RSI

[LC10] Ziyang Liu and Yi Chen. Return specification inference and result clustering for keyword search on XML. ACM Transactions on Database Systems, 35(2):10:1– 10:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2010:IXS

[LHC10] Ziyang Liu, Yu Huang, and Yi Chen. Improving XML search by generating and utilizing informative result snippets. ACM Transactions on Database Systems, 35(3):19:1–19:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lee:2010:SCE

[LWHK10] Ki-Hoon Lee, Kyu-Young Whang, Wook-Shin Han, and Min-Soo Kim. Structural consistency: enabling XML keyword search to eliminate spurious results consistently. VLDB Journal: Very Large Data Bases, 19(4):503–529, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lin:2010:NXK

[LWXZ10] Xudong Lin, Ning Wang, De Xu, and Xiaoning Zeng. A novel XML keyword query approach using entity subtree. The Journal of Systems and Software, 83(6): 990–1003, June 2010. CODEN JS-SODM. ISSN 0164-1212.

Martens:2010:CDP

[MNS10] Wim Martens, Frank Neven, and Thomas Schwentick. Complexity of decision problems for XML schemas and chain regular expressions. SIAM Journal on Computing, 39(4):1486–1530, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miori:2010:DTI

[MRA10] Vittorio Miori, Dario Russo, and Massimo Aliberti. Domotic technologies incompatibility becomes user transparent. Communications of the ACM, 53(1):

153–157, January 2010. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pilgrim:2010:HR

[Pil10] Mark Pilgrim. HTML5: up and running. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-80602-7. 250 pp. LCCN ????

Pozrikidis:2013:XSC

[Poz13] C. Pozrikidis. XML in scientific computing. Chapman and Hall/CRC numerical analysis and scientific computing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-1227-X (hardback). xv + 243 pages pp. LCCN Q183.9 .P69 2013.

Romei:2010:XDM

[RT10] Andrea Romei and Franco Turini. XML data mining. Software— Practice and Experience, 40(2): 101–130, February ??, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schmitt:2010:CC

[Sch10] Christopher Schmitt. CSS cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 0-596-15593-X (paperback). xxiv + 702 pp. LCCN TK5105.888 .S3524 2010. Foreword by Dan Cederholm.

Stamerjohanns:2010:TLC

[SKG^{+}10] Heinrich Stamerjohanns, Michael Kohlhase, Deyan Ginev, Catalin David, and Bruce Miller. Transforming large collections of scientific publications to XML. Mathematics in Computer Science, 3 (3):299–307, May 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://www.springerlink.com/
openurl.asp?genre=article&
issn=1661-8270&volume=3&issue=
3&spage=299.

Silvasti:2010:ELX

[SSSS10] P. Silvasti, S. Sippu, and E. Soisalon-Soininen. Evaluating linear XPath expressions by pattern-matching automata. J.UCS: Journal of Universal Computer Science, 16(5):833– ??, ???? 2010. CODEN ???? ISSN 0948-6968. URL

http://www.jucs.org/jucs_16_
5/evaluating_linear_xpath_
expressions.

Stark:2010:BIA

[Sta10] Jonathan Stark. Building iPhone apps with HTML, CSS, and JavaScript: Making App Store apps without Objective-C or Cocoa. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 1-4493-8023-9, 0-596-80578-0. xv + 166 pp. LCCN ????

Tagarelli:2010:SCX

[TG10] Andrea Tagarelli and Sergio Greco. Semantic clustering of XML documents. ACM Transactions on Information Systems, 28 (1):3:1–3:??, January 2010. CO-DEN ATISET. ISSN 1046-8188.

TenCate:2010:TCL

[TS10] Balder Ten Cate and Luc Segoufin. Transitive closure logic, and nested tree walking automata, and Xpath. Journal of the ACM, 57(3):18:1–18:??, March 2010. CODEN JACOAH. ISSN 0004-5411.

Vaughan-Nichols:2010:WHR

[VN10] Steven J. Vaughan-Nichols. Will HTML 5 restandardize the Web? Computer, 43(4):13–15, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).